what is the answer to this question.
ayie is 8 year older than icas. the sum of their ages is 32. how old is each now.
pls show the solution thank you!!!

Answer by ptaylor(1894)

Let x=icas age now
then 8+x=ayie's age now
We are told that the sum of their ages is 32
So,
x+8+x=32 subtract 8 from each side
x+8-8+x=32-8 collect like terms
2x=24 divide each side by 2
x=12------------------icas age now
8+x=8+12=20------------------ayie's age now
CK
20+12=32
32=32
Hope this helps---ptaylor